# Generalizable Spectral Embedding with an Application to UMAP

### v1

@@ 5.2 Scalability
Noteworthy, GrEASE not only generalizes effectively but also does so more quickly than the most scalable (yet non-generalizable) existing methods. Fig. 2(b) demonstrates this point on the toy moon dataset - a 2D moon linearly embedded into 10D input space. To evaluate scalability, we measured the computation time required for SE approximation, for an increasing number of samples. We compared the results with the three most popular methods for sparse matrix decomposition, which are currently the fastest implementations: ARPACK (Lehoucq et al., 1998), LOBPCG (Benner and Mach, 2011), and AMG (Brandt et al., 1984). For each number of samples, we calculated the Laplacian matrix that is 99% sparse. Each method was executed five times, initialized with different seeds. As discussed in Sec. 4, GrEASE demonstrates approximately linear time complexity, and indeed, for higher numbers of samples, GrEASE converges significantly faster.
@@ 5.3 NUMAP - generalizable UMAP
In this section, we demonstrate NUMAP’s ability to preserve global structure, while enabling fast inference of test points, and it’s ability to enable time-series UMAP visualization. We compare our results with P. UMAP and NUMAP-SN (NUMAP architecture using SpectralNet instead of GrEASE). We consider four real-world datasets: CIFAR10 (via their CLIP embedding); Appliances Energy Prediction dataset; Wine (Aeberhard and Forina, 1992); and Banknote Authentication (Lohweg, 2012). For additional technical details regarding the datasets, architectures and training procedures, we refer the reader to Appendix G.
@@
Tab. 2 presents our results on the real-world datasets. The local (i.e., kNN) results are comparable across the three methods. However, NUMAP consistently better captures the global structure (based on the lower GS). In other words, NUMAP achieves comparable local preservation results with P. UMAP, while possessing more global structure expressivity. Also, the table shows that GrEASE is necessary to achieve these results, which are not reproduced with SpectralNet.
In Fig. 1, we supplement the empirical results with qualitative examples. Fig. 1a presents three simple non-linear synthetic 3-dimensional structures and their 2-dimensional visualizations using UMAP (non-parametric), P. UMAP and NUMAP. UMAP (using its default configuration, SE initialization) accurately preserves the global structure in its 2-dimensional representations, but lack the ability to generalize to unseen points. Among the generalizable methods (i.e., P. UMAP and NUMAP), P. UMAP fails to preserve the global structure: in the top two rows, it does not separate the clusters, while in the bottom row, it introduces undesired color overlaps. In contrast, NUMAP effectively preserves these separations and avoids the unnecessary overlapping. These examples are particularly insightful, as P. UMAP fails to visualize correctly even these simple datasets. Fig. 1b further demonstrates NUMAP’s superior ability to preserve global structure, as evidenced by the improved class separation in the Banknote dataset.
Table 2: NUMAP preserves global structure of real-world datasets. A comparison between NUMAP and P. UMAP visualization on real-world datasets. The values are the mean and standard deviation of the measures on the test set, over 5 runs. NUMAP is superior in preserving global structure.
@@
kNN $\uparrow$ P. UMAP 0.886±0.043 - $0.922_{\pm 0.027}$ 0.927±0.031
NUMAP-SN 0.862±0.008 - $0.956_{\pm 0.028}$ 0.975±0.022
NUMAP (ours) 0.874±0.023 - 0.956±0.033 0.988±0.002
GS $\downarrow$ P. UMAP 0.102±0.043 0.769±0.262 0.502±0.146 0.685±0.035
NUMAP-SN 0.460±0.267 0.255±0.044 0.617±0.193 0.815±0.168
NUMAP (ours) 0.089±0.042 0.244±0.015 0.461±0.161 0.570±0.122
Time-series visualization.
Fig. 6 shows a simulation time-series data, which can be viewed as a simulation of cellular differentiation. Specifically, we may consider differentiation of hematopoietic stem cells (also known as blood stem cells), which are known to differentiate into many types of blood cells, to T-cells. The process involves two kinds of cells (represented by their gene expressions; red and blue samples in the figure). One represents stem cells, while the other T-cells. A group of cells (colored in pink in the figure) then gradually transitions from stem cells to T-cells. At the top row we use UMAP to visualize each time step, while at the bottom we train NUMAP on the first two time-steps and only inference the rest. UMAP is inconsistent over time-steps, which makes it impractical for understanding change and progression. It also has to train the embeddings each time-step separately. In contrast, NUMAP only trains on the first two time-steps and the embeddings of the later time-steps are immediate from inference. This also enables consistency over time, and makes the trend and process visible and understandable.
Figure 6: Time-series visualization using NUMAP. Visualization of a dynamical system using UMAP and NUMAP. NUMAP is both consistent and does not require training after the first two time-steps.
@@ 6 Conclusions
We first introduced GrEASE, a deep-learning approach for approximate SE. GrEASE addresses the three primary drawbacks of current SE implementation: generalizability, scalability and eigenvectors separation. By incorporating a post-processing diagonalization step, GrEASE enables eigenvectors separation without compromising generalizability or scalability. Remarkably, this one-shoot post-processing operation lays the groundwork for a wide range of new applications of SE, which would not have been possible without its scalable and generalizable implementation. It also presents a promising pathway to enhance current applications of SE.
In particular, we presented NUMAP, a novel application of GrEASE for generalizable UMAP visualization. We believe the integration of SE with deep learning can have a significant impact on unsupervised learning methods. Further research should delve into exploring the applications of SE across various fields.
@@ Appendix A Proof of Lemma 1
The Rayleigh Quotient of a positive semi-definite matrix $L\in\mathbb{R}^{n\times n}$ with eigenvectors $v_{1},\dots,v_{n}$ corresponding to the eigenvalues $\lambda_{1}\leq\cdots\leq\lambda_{n}$, $R_{L}$ satisfies $arg\min_{||v||=1}R_{L}(v)=v_{1}$ and for each $i>1$ $arg\min_{||v||=1}R_{L}(v)=v_{i}$ for $v\perp v_{1},\dots,v_{i-1}$ (Li, 2015).
@@
As $V$ minimizes $R_{L}$, we get that $min_{U}R_{L}(U)=R_{L}(V)=\sum_{i=1}^{k}\lambda_{i}$, where $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ are the eigenvalues of $L$. This yields
$R_{L}(A)=\text{Tr}(A^{T}LA)=\sum_{i=1}^{k}\lambda_{i}$
@@ 6       Forward propagate $X$ and compute inputs to orthogonalization layer $\tilde{Y}$
7       Compute the $QR$ factorization $QR=\tilde{Y}$
@@ 12       Forward propagate $x_{1},\dots,x_{m}$ to get $y_{1},\dots,y_{m}$
3 Forward propagate all $X_{i}$ and obtain $F_{\theta}$ outputs $Y_{i}\in\mathbb{R}^{m\times k+1}$
@@ C.1 Time and Space Complexity
Specifying the exact complexity of the method is difficult, As this is a non-convex optimization problem, However, we can discuss the following approximate complexity analysis. Assuming constant input and output dimensions and a given network architecture, we can take a general view on the complexity of each iteration by the batch size $m$. The heaviest computational operations at each iteration are the nearest-neighbors search, the QR decomposition and the loss computation (i.e., computation of the Rayleigh Quotient). For the nearest-neighbor search, we can use approximation techniques (e.g, LSH Gionis et al. (1999)) which work in almost linear complexity by $m$. A naive implementation of the QR decomposition would lead to an $\mathcal{O}(m^{2})$ time complexity. The loss computation also takes $\mathcal{O}(m^{2})$ due to the required matrix multiplication. Thereby, the complexity of each iteration is quadratic by the batch size. This is comparable to other approximation techniques such as LOBPCG Benner and Mach (2011) (which also utilizes sparse matrix operations techniques for faster implementation). However, GrEASE leverages stochastic training, allowing each iteration to consider only a batch of the data, rather than the entire dataset.
@@ C.2 Graph Construction
To best capture the structure of the input manifold $\mathcal{D}$, given by a finite number of samples $\mathcal{X}$, we use a similar graph construction method used by Gomez et al. in UMAP (McInnes et al., 2018), proven to capture the local topology of the manifold at each point. However, as opposed to the method in (McInnes et al., 2018), GrEASE does not compute the graph of all points, which can lead to scalability hurdles and impose significant memory demands. Instead, GrEASE either computes small graphs on each batch, or can be provided by the user with an affinity matrix $W$ corresponding to $\mathcal{X}$. Our practical construction of the graph affinity matrix $W$ is as follows:
@@ D.1 Formalization of GS
First, we remind Grassmann distance (see Def. 1). Grassmann distance is a metric function between equidimensional linear subspaces, where each is represented by an orthogonal matrix containing the basis as its columns. In other words, this is a metric which is invariant under multiplication by an orthogonal matrix.
@@
$t$ is a hyper-parameter of GS. A reasonable choice would be to take $t=2$, which is equivalent to measure the Grassmann distance between the Fiedler vectors of the Laplacians. The Fiedler vector is known for its hold of the most important global properties. The larger $t$, the more complicated structures are taken into consideration in the GS computation (which is not necessary desired).
@@ D.2 Additional GS examples
Fig. 7 depicts two additional demonstrations of the alignment between the intuitive expectation and the GS results on two toy dataset. The basic global structure of both of these datasets is two distinct clusters. This structure is indeed captured by GS. However, kNN gives perfect score also when the one of the clusters is separated. Silhouette score favorites the 2-points embedding. Namely, it trade-offs local structure (i.e., giving lower score for preserving local structure, even when the global properties are the same).
Appendix E Additional results
The full results of Fig. 5 are summarized in Tab. 3.
Table 3: A comparison between GrEASE and SpectralNet dimensional SE and Fiedler Vector (FV) approximation on real-world datasets. The values are the mean and standard deviation of the $sin^{2}$ distance between the predicted and true eigenvector, over 10 runs. Lower is better. GrEASE ability to separate the eigenvectors is evident.
@@ Appendix E Additional results
Appliances GrEASE $0.063_{\pm 0.002}$ $0.094_{\pm 0.007}$ $0.109_{\pm 0.001}$
SpectralNet $0.307_{\pm 0.047}$ $0.530_{\pm 0.114}$ $0.401_{\pm 0.106}$
KMNIST GrEASE $0.0.044_{\pm 0.002}$ $0.101_{\pm 0.010}$
SpectralNet $0.372_{\pm 0.174}$ $0.396_{\pm 0.137}$
Parkinsons GrEASE $0.056_{\pm 0.006}$
SpectralNet $0.229_{\pm 0.138}$
Appendix F Fine-Tuning GrEASE with UMAP loss
One way to get a generalizable version of UMAP may be an extension of GrEASE by fine-tuning the network with UMAP loss. We tried that idea, but were forced to stop this direction, as we stumbled upon the well-known catastrophic forgetting case.
Figure 8 presents an experiment on the simple 2circles dataset. Each row is represented the same experiment, run with a different seed. We trained GrEASE to output the 2D SE of the 2circles dataset, as shown in the left column. Then, we initialized a new network, with the same architecture, with the pre-trained weights from GrEASE. This network was trained with UMAP loss, as in (Sainburg et al., 2021). We tried different learning-rates for fine-tuning, to best match the desired UMAP embedding (i.e. retaining the local structure), without losing the global structure (e.g., separation of the two clusters). Unfortunatly, there was no learning-rate that matched our goals.
Figure 8: The catastrophic forgetting phenomenon when fine-tuning GrEASE to much UMAP performance on the 2circles dataset. Each column represents a fine-tuning using a different learning-rate. Each row is a repetition, initialized with a different seed.

### v2

@@ 5.2 Scalability
Noteworthy, Sep-SpectralNet not only generalizes effectively but also does so more quickly than the most scalable (yet non-generalizable) existing methods. Fig. 2(b) demonstrates this point on the toy moon dataset - a 2D moon linearly embedded into 10D input space. To evaluate scalability, we measured the computation time required for SE approximation, for an increasing number of samples. We compared the results with the three most popular methods for sparse matrix decomposition, which are currently the fastest implementations: ARPACK (Lehoucq et al., 1998), LOBPCG (Benner & Mach, 2011), and AMG (Brandt et al., 1984). For each number of samples, we calculated the Laplacian matrix that is 99% sparse. Each method was executed five times, initialized with different seeds. Notably, for higher numbers of samples, Sep-SpectralNet converges significantly faster. Notably, Sep-SpectralNetmaintains the same scalability as SpectralNet, since the separation step incurs negligible overhead compared to the overall training time.
@@ 5.3 NUMAP - generalizable UMAP
In this section, we demonstrate NUMAP’s ability to preserve global structure, while enabling fast inference of test points, and it’s ability to enable time-series UMAP visualization. We compare our results with P. UMAP, with the target dimensionality set to 2. We refer to App. E for various ablations.
We consider four real-world datasets: CIFAR10 (via their CLIP embedding); Appliances Energy Prediction dataset; Wine (Aeberhard & Forina, 1992); Banknote Authentication (Lohweg, 2012). For additional technical details regarding the datasets, architectures and training procedures, we refer to App. G.
@@
Local Structure.
Tab. 2 presents the kNN results on the real-world datasets. These results are comparable across the methods. Namely, NUMAP does not compromise on local structure.
Figure 6: NUMAP and P. UMAP visualization of the Banknote dataset test set. NUMAP’s enhanced ability to preserve global structure is evident in the clearer separation between clusters compared to P. UMAP.
Global Structure.
Tab. 3 presents the GS results on these datasets. The results underscore that NUMAP better captures the global structure. Namely, NUMAP enhances global structure preservation while not compromising local structure preservation.
In Fig. 1 and Fig. 6, we supplement the empirical results with qualitative examples. Fig. 1 presents simple non-linear synthetic 3-dimensional structures and their 2-dimensional visualizations using UMAP (non-parametric), P. UMAP and NUMAP. These examples are particularly insightful, as the expected outcome of a good visualization is known. UMAP (using its default configuration, SE initialization) accurately preserves the global structure in its 2-dimensional representations, but lack the ability to generalize to unseen points. Among the generalizable methods (Fig. 1b), P. UMAP fails to preserve the global structure: in the top two rows, it does not separate the clusters, while in the bottom row, it introduces undesired color overlaps. In contrast, NUMAP effectively preserves these separations and avoids unnecessary overlapping. Fig. 6 further demonstrates NUMAP’s ability to preserve global structure, as evidenced by the improved class separation in the Banknote dataset.
@@
kNN $\uparrow$ P. UMAP 0.908±0.004 - 0.953±0.033 0.927±0.023
NUMAP (ours) 0.905±0.004 - 0.950±0.024 0.986±0.004
Table 2: NUMAP local structure preservation is comparable with P. UMAP. kNN accuracy results of P. UMAP and NUMAP on labeled real-world datasets (Appliances dataset is excluded due to the absence of labels). The values are the mean and standard deviation measures on the test set, over 10 runs.
Metric Method Cifar10 Appliances Wine Banknote
GS $\downarrow$ P. UMAP 0.133±0.069 0.710±0.293 0.549±0.180 0.722±0.079
NUMAP (ours) 0.054±0.021 0.261±0.020 0.429±0.124 0.618±0.131
p-value 0.004 0.0002 0.118 0.056
Table 3: NUMAP better preserves global structure. P. UMAP and NUMAP Grassmann Score (GS) results on real-world datasets. The values are the mean and standard deviation measures on the test set, over 10 runs. The best in mean is highlighted. Notably, NUMAP achieves better global structure preservation.
Time-series data visualization.
Fig. 7 shows a simulation time-series data, which can be viewed as a simulation of cellular differentiation. Specifically, we may consider differentiation of hematopoietic stem cells (also known as blood stem cells), which are known to differentiate into many types of blood cells, to T-cells. The process involves two kinds of cells (represented by their gene expressions; red and blue samples in the figure). One represents stem cells, while the other T-cells. A group of cells (colored in pink in the figure) then gradually transitions from stem cells to T-cells. At the top row we use UMAP to visualize each time step, while at the bottom we train NUMAP on the first two time-steps and only inference the rest. UMAP is inconsistent over time-steps, which makes it impractical for understanding change and progression. It also has to train the embeddings at each time-step separately. In contrast, NUMAP only trains on the first two time-steps, and the embeddings of the later time-steps are immediate from inference. This also enables consistency over time, and makes the trend and process visible and understandable.
Figure 7: Time-series data visualization using NUMAP. A comparison between UMAP and NUMAP visualizations of a dynamical system. In the UMAP visualization, the structure and cluster positions vary significantly across timesteps, hindering the interpretability of the underlying process. In contrast, NUMAP produces consistent structures over time, clearly revealing the transition of the pink points between the red and blue regions. Notably, NUMAP requires training only on the first two timesteps, enabling efficient generalization to the rest of the sequence.
@@ 6 Conclusions
We introduced Sep-SpectralNet, a deep-learning approach for approximate SE. Sep-SpectralNet addresses the three primary drawbacks of current SE implementation: generalizability, scalability and eigenvectors separation. By incorporating a post-processing diagonalization step, Sep-SpectralNet enables eigenvectors separation without compromising SpectralNet’s generalizability or scalability. Remarkably, this one-shot post-processing operation lays the groundwork for a wide range of new applications of SE, which would not have been possible without its scalable and generalizable implementation. It also presents a promising pathway to enhance current applications of SE.
In particular, we also presented NUMAP, a novel extension of Sep-SpectralNet for generalizable UMAP visualization. We believe the integration of generalizable SE with deep learning can have a significant impact on unsupervised learning methods. However, exploring alternative inputs to the UMAP loss network is an interesting direction for future work. Finally, further research should delve into exploring the applications of SE across various fields.
@@ Appendix A Proof of Lemma 1
[…]
@@
As $V$ minimizes $R_{L}$, we get that $\min_{U}R_{L}(U)=R_{L}(V)=\sum_{i=1}^{k}\lambda_{i}$, where $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ are the eigenvalues of $L$. This yields
$R_{L}(A)=\operatorname{Tr}(A^{T}LA)=\sum_{i=1}^{k}\lambda_{i}$
@@ 6   Forward propagate $X$ and compute inputs to orthogonalization layer $\tilde{Y}$
7   Compute the QRQR factorization $QR=\tilde{Y}$
@@ 12   Forward propagate $x_{1},\dots,x_{m}$ to get $y_{1},\dots,y_{m}$
3 Forward propogate all $X_{i}$ and obtain $F_{\theta}$ outputs $Y_{i}\in\mathbb{R}^{m\times k+1}$
@@ C.1 Time and Space Complexity
Specifying the exact complexity of the method is difficult, As this is a non-convex optimization problem, However, we can discuss the following approximate complexity analysis. Assuming constant input and output dimensions and a given network architecture, we can take a general view on the complexity of each iteration by the batch size $m$. The heaviest computational operations at each iteration are the nearest-neighbors search, the QR decomposition and the loss computation (i.e., computation of the Rayleigh Quotient). For the nearest-neighbor search, we can use approximation techniques (e.g, LSH Gionis et al. (1999)) which work in almost linear complexity by $m$. A naive implementation of the QR decomposition would lead to an $\mathcal{O}(m^{2})$ time complexity. The loss computation also takes $\mathcal{O}(m^{2})$ due to the required matrix multiplication. Thereby, the complexity of each iteration is quadratic by the batch size. This is comparable to other approximation techniques such as LOBPCG Benner & Mach (2011) (which also utilizes sparse matrix operations techniques for faster implementation). However, Sep-SpectralNet leverages stochastic training, allowing each iteration to consider only a batch of the data, rather than the entire dataset.
@@ C.2 Graph Construction
To best capture the structure of the input manifold $\mathcal{D}$, given by a finite number of samples $\mathcal{X}$, we use a similar graph construction method used by Gomez et al. in UMAP (McInnes et al., 2018), proven to capture the local topology of the manifold at each point. However, as opposed to the method in (McInnes et al., 2018), Sep-SpectralNet does not compute the graph of all points, which can lead to scalability hurdles and impose significant memory demands. Instead, Sep-SpectralNet either computes small graphs on each batch, or can be provided by the user with an affinity matrix $W$ corresponding to $\mathcal{X}$. Our practical construction of the graph affinity matrix $W$ is as follows:
@@ D.1 Formalization of GS
Grassmann distance (see Def. 1) is a metric function between equidimensional linear subspaces, where each is represented by an orthogonal matrix containing the basis as its columns. In other words, this is a metric which is invariant under multiplication by an orthogonal matrix.
@@
$t$ is a hyper-parameter of GS. A reasonable choice would be to take $t=2$, which is equivalent to measure the Grassmann distance between the Fiedler vectors of the Laplacians. The Fiedler vector is known for its hold of the most important global properties. The larger $t$, the more complicated structures are taken into consideration in the GS computation (which is not necceray desired).
@@ D.2 Additional GS examples
Fig. 8 depicts two additional demonstrations of the alignment between the intuitive expectation and the GS results on two toy dataset. The basic global structure of both of these datasets is two distinct clusters. This structure is indeed captured by GS. However, kNN gives perfect score also when the one of the clusters is separated. Silhouette score favourites the 2-points embedding. Namely, it trade-offs local structure (i.e., giving lower score for preserving local structure, even when the global properties are the same).
Appendix E NUMAP Ablation Study
In this section, we start by considering two ablations of NUMAP, which for convenience we name NUMAP-SN and NUMAP-FT. NUMAP-SN is a replication of NUMAP’s pipeline with SpectralNet replacing Sep-SpectralNet. That is, we show that eigenvector separation is necessary for NUMAP.
As for NUMAP-FT, this name refers to another method to get a generalizable version of UMAP by an extension of Sep-SpectralNet  by fine-tuning the network with UMAP loss. We tried that idea, but were forced to stop this direction, as we stumbled upon the well-known catastrophic forgetting case.
Tab. 4 extends Tab. 2 with the results of NUMAP-SN and NUMAP-FT. Notably, NUMAP-SN local structure preservation are comparable with NUMAP and P. UMAP, as expected, while NUMAP-FT consistently fails in local structure preservation.
Tab. 5 extends Tab. 3 with the results of NUMAP-SN and NUMAP-FT. Notably, NUMAP-SN is less consistent than NUMAP in global structure preservation, resulting in worse performance on the Cifar10 and Banknote datasets. NUMAP-FT fails in global structure preservation on the Cifar10 and Appliances datasets. Although it achieves better GS on the Wine dataset, and comparable global structure preservation on the Banknote dataset, it compromises a lot on local structure preservation (see Tab. 4).
Figure 9 presents an experiment on the simple 2circles dataset. Each row is represented the same experiment, run with a different seed. We trained Sep-SpectralNet to output the 2D SE of the 2circles dataset, as shown in the left column. Then, we initialized a new network, with the same architecture, with the pre-trained weights from Sep-SpectralNet. This network was trained with UMAP loss, as in (Sainburg et al., 2021). We tried different learning-rates for fine-tuning, to best match the desired UMAP embedding (i.e. retaining the local structure), without losing the global structure (e.g., separation of the two clusters). Unfortunately, there was no learning-rate that matched our goals.
Metric Method Cifar10 Appliances Wine Banknote
kNN $\uparrow$ P. UMAP 0.908±0.004 - 0.953±0.033 0.927±0.023
NUMAP (ours) 0.905±0.004 - 0.950±0.024 0.986±0.004
NUMAP-SN 0.903±0.002 - 0.956±0.028 0.963±0.032
NUMAP-FT 0.577±0.081 - 0.364±0.053 0.686±0.046
Table 4: Local structure preservation ablation. An extension of Tab. 2 with NUMAP-SN and NUMAP-FT.
Metric Method Cifar10 Appliances Wine Banknote
GS $\downarrow$ P. UMAP 0.133±0.069 0.710±0.293 0.549±0.180 0.722±0.079
NUMAP (ours) 0.054±0.021 0.261±0.020 0.429±0.124 0.618±0.131
NUMAP-SN 0.281±0.357 0.262±0.038 0.441±0.182 0.780±0.197
NUMAP-FT 0.680±0.334 0.348±0.292 0.002±0.001 0.635±0.141
Table 5: Global structure preservation ablation. An extension of Tab. 3 with NUMAP-SN and NUMAP-FT.
Figure 9: The catastrophic forgetting phenomenon when fine-tuning Sep-SpectralNet to match UMAP performance on the 2circles dataset. Each column represents a fine-tuning using a different learning-rate. Each row is a repetition, initialized with a different seed.
Another important question is the necessity of the residual connections. For that, we conduct another ablation study on the Cifar10 dataset. The results are shown in Tab. 6 and Tab. 7. Notably, while the skip connections does not improve local structure preservation, they are crucial for global structure preservation, indicated by the lower GS.
Metric Method Cifar10
kNN $\uparrow$ NUMAP (no residual connections) 0.907±0.003
NUMAP 0.905±0.004
Table 6: Ablation on residual connections for local structure preservation shows comparable performance with and without their use.
Metric Method Cifar10
GS $\downarrow$ NUMAP (no residual connections) 0.191±0.293
NUMAP 0.054±0.021
Table 7: Ablation on residual connections for global structure preservation shows improved performance when they are used.
Appendix F Additional results
The full results of Fig. 5 are summarized in Tab. 8. The visualizations corresponding to Tab. 3 are depicted in Fig. 10. Fig. 11 extends Fig. 7 with P. UMAP.
Tab. 9 extends Tab. 4 with additional two datasets: MNIST (Deng, 2012) and FashionMNIST (Xiao et al., 2017). Tab. 10 correspondingly extends Tab. 5.
Table 8: A comparison between Sep-SpectralNet and SpectralNet dimensional SE and Fiedler Vector (FV) approximation on real-world datasets. The values are the mean and standard deviation of the $\text{sin}^{2}$ distance between the predicted and true eigenvector, over 10 runs. Lower is better. Sep-SpectralNet ability to separate the eigenvectors is evident.
@@ Appendix F Additional results
Appliances Sep-SpectralNet $063_{\pm 0.002} 0.0$$094_{\pm 0.007}$ $0.109_{\pm 0.001}$ -
SpectralNet $0.307_{\pm 0.047}$ $0.530_{\pm 0.114}$ $0.401_{\pm 0.106}$ -
KMNIST Sep-SpectralNet $0.0.044_{\pm 0.002}$ $0.101_{\pm 0.010}$ - -
SpectralNet $0.372_{\pm 0.174}$ $0.396_{\pm 0.137}$ - -
Parkinsons Sep-SpectralNet $0.056_{\pm 0.006}$ - - -
SpectralNet $0.229_{\pm 0.138}$ - - -
Figure 10: Test set Visualizations of NUMAP, NUMAP-SN and P. UMAP on the datasets corresponding to Tab. 2 and Tab. 3.
Figure 11: Fig. 7 extension with P. UMAP.
Metric Method Cifar10 Appliances Wine Banknote Mnist FashionMnist
kNN $\uparrow$ P. UMAP 0.908±0.004 - 0.953±0.033 0.927±0.023 0.801±0.010 0.717±0.006
NUMAP (ours) 0.905±0.004 - 0.950±0.024 0.986±0.004 0.758±0.009 0.695±0.003
NUMAP-SN 0.903±0.002 - 0.956±0.028 0.963±0.032 0.750±0.010 0.695±0.006
NUMAP-FT 0.577±0.081 - 0.364±0.053 0.686±0.046 0.329±0.103 0.153±0.035
Table 9: An extension of Tab. 4 with the MNIST and FashionMNIST datasets. Local preservation results are corresponding to Tab. 4. Namely, NUMAP, P. UMAP and NUMAP-SN are comparable, with NUMAP-FT failing to preserve local structure.
Metric Method Cifar10 Appliances Wine Banknote Mnist FashionMnist
GS $\downarrow$ P. UMAP 0.133±0.069 0.710±0.293 0.549±0.180 0.722±0.079 0.311±0.059 0.029±0.006
NUMAP (ours) 0.054±0.021 0.261±0.020 0.429±0.124 0.618±0.131 0.304±0.042 0.032±0.022
NUMAP-SN 0.281±0.357 0.262±0.038 0.441±0.182 0.780±0.197 0.405±0.199 0.036±0.016
NUMAP-FT 0.680±0.334 0.348±0.292 0.002±0.001 0.635±0.141 0.411±0.153 0.272±0.146
Table 10: An extension of Tab. 5 with the MNIST and FashionMNIST datasets. Global preservation results on MNIST are comparable between UMAP and P. UMAP, while superior to NUMAP-SN and NUMAP-FT. Regarding FashionMNIST, NUMAP, FashionMNIST, and NUMAP-SN are comparable, with NUMAP-FT having worse preservation.